OpenStudy (ammarah):

Find the discriminant and use it to determine if the solution has one real, two real, or two imaginary solutions. 1. x^ +4x+3=0 2. x^ -2x +1=0 3. -x^ -x =4

OpenStudy (anonymous):

I will do #1, you'll do the rest. x^2 + 4x + 3 = 0 The discriminant is the value of b^2 - 4ac a= 1, b = 4, c = 3 so b^2 - 4ac = 16 - 4(1)(3) = 4 So the discrimiant is 4. If the discrmiant is greater than zero and a perfect square...like 4 in our problem, then the roots are real and rational. Done. One note...if the discriminant is less than zero, like -12, the roots are imaginary If the the discriminant is zero, roots are real and equal. If the discrminnant is greater than zero but NOT a perfect square, then the roots are real, unequal, and irrational.

OpenStudy (anonymous):

Now, you should be able to do #2 and #3 or any problem.